# Meiying Chen and Zhiyao Duan

### v5

@@ 1 Introduction
To our best knowledge, ControlVC is the first neural VC system that achieves time-varying controls on pitch and speed. ControlVC performs a generic control on pitch and speed but not trying to mimic those of the target speaker. It's worth noting that works such as [20, 21, 22] may have the potential to achieve time-varying speed and/or pitch control, but substantial modifications of the algorithms and necessary experiments would be needed to support such claim, which is not the primary focus of their research. To validate the proposed system, we conduct extensive subjective and objective evaluations and compare it with two self-constructed baselines as no existing systems were available. Experimental results show that ControlVC realizes a good level of time-varying controllability on pitch and speed, while achieving significantly better naturalness and timbre similarity than the comparison methods.
@@ 2.6 HiFi-GAN Neural Vocoder
To construct the encoded discrete representation, the linguistic and pitch embeddings are up-sampled, and the utterance-level speaker embedding is copied to match the same frame rate. These three embeddings are then concatenated into the intermediate representation $\mathbf{z}=(\mathbf{z}^{(p)},\mathbf{z}^{(l)},\mathbf{z}^{(s)})$, which is then fed into a neural vocoder to generate the waveform. We use HiFi-GAN's [19] vocoder but modify the original implementation so that it directly accepts the discrete and continuous mixed representation $\mathbf{z}$ as input. The generator includes a set of transposed convolution blocks to increase sample rate of the discrete representation and a residual block with dilated layers to increase the receptive filed. The discriminator contains two types of sub-discriminators: five multi-period discriminators (MPD) and three multi-scale discriminators (MSD).
@@
$L_{G}$ $=\sum_{k=1}^{K}\biggl{[}L_{Adv}(G;D_{k})+\lambda_{fm}L_{FM}(G;D_{k})\biggr{]}$ (1)
@@
$L_{Mel}(G)$ $=\mathbb{E}_{x,\hat{x}}\biggl{[}||\phi(x)-\phi(G(x))||_{1}\biggr{]},$ (4)
@@ 3.4 Experimental Setup
The VC experiments are performed among all 90 pairs of 10 test speakers. Each utterance of one test speaker is converted to each of the other 9 speakers' voices. Each test speaker reads a different set of sentences. All the 10 speakers and their sentences are unseen during training. Therefore, the conversion is non-parallel and zero-shot.
In this experiment, we apply control curves for speed and/or pitch. Four control settings are tested: ``No Control" - traditional voice conversion without any explicit control; ``Pitch Only" and ``Speed Only" denote pitch or speed control but not both; ``Speed+Pitch" means both aspects are controlled. We test two curves for pitch control: stressing (i.e., pitch rising abruptly then going down gradually) and rising, and three curves for speed control: parabola, speed up and slow down. The control settings and the control curves are drawn with equal probability for each conversion. We perform both subjective and objective evaluations to assess the conversion quality, intelligibility and controllability of the proposed system.
@@ 3.6 Subjective Evaluation
Audio Quality Test. In the first test, we use mean opinion score (MOS) [35] to assess the naturalness and timbre similarity of the converted speech. Study participants are presented with a set of utterances including one source, one target and several converted utterances. Each set is referred to as a sample. The converted utterances from ControlVC and the two baselines are presented in random order. For each sample, participants are asked to rate between 1-5 on the naturalness of the converted utterances and the timbre similarity between the target and the converted utterance. Higher scores are better. Each participant is asked to complete at least 6 samples, i.e., 36 ratings for 18 converted utterances from 3 comparison methods. Participants are allowed to complete more samples. In total, 233 samples are evaluated, resulting in 1398 ratings. As shown in Fig. 2, ControlVC archives the best MOS among three comparison methods in all control settings. In addition, comparing the three with-control settings with ``No Control'', we see that applying controls only slightly decreases the speech quality of the converted speech from ControlVC.
@@
Controllability Test. The second test assesses the controllability of the proposed ControlVC method. We do not include the two baselines in this test due to their poor audio quality in the previous test. The participants are presented with uncontrolled and controlled conversion results, along with a figure of the corresponding control curve(s). The participants are then asked to assess how accurately the curve describes the change of pitch or speed between the uncontrolled and controlled conversions on a scale of 1 to 5, with 1 being ``not at all accurate", 3 being``moderately accurate" and 5 being ``very accurate". Each participant is asked to complete 6 rounds of tests, with each round containing a pitch control, a speed control and a pitch+speed control of the same source-target reference pair. Participants are allowed to complete more rounds. In total, participants completed 169 rounds of tests, resulting in 676 ratings.
@@ 4 Conclusions
In this paper, we proposed a controllable voice conversion system named ControlVC, which allows users to impose time-varying controls on pitch and speed in voice conversion. The converted utterance maintains the source utterance's linguistic content, mimics the target speaker's timbre, and sounds natural while following the user input pitch and/or speed controls. Both subjective and objective evaluation results suggest that ControlVC is able to perform pitch and speed control while producing high-quality conversions.

### v6

@@ 1 Introduction
To our best knowledge, ControlVC is the first neural VC system that achieves time-varying controls on pitch and speed. ControlVC performs a generic control on pitch and speed but not trying to mimic those of the target speaker. It’s worth noting that works such as [20, 21, 22] may have the potential to achieve time-varying speed and/or pitch control, but substantial modifications of the algorithms and necessary experiments would be needed to support such claim, which is not the primary focus of their research. To validate the proposed system, we conduct extensive subjective and objective evaluations and compare it with two self-constructed baselines as no existing systems were available. Experimental results show that ControlVC realizes a good level of time-varying controllability on pitch and speed, while achieving significantly better naturalness and timbre similarity than the comparison methods.
@@ 2.6 HiFi-GAN Neural Vocoder
To construct the encoded discrete representation, the linguistic and pitch embeddings are up-sampled, and the utterance-level speaker embedding is copied to match the same frame rate. These three embeddings are then concatenated into the intermediate representation $\mathbf{z}=(\mathbf{z}^{(p)},\mathbf{z}^{(l)},\mathbf{z}^{(s)})$, which is then fed into a neural vocoder to generate the waveform. We use HiFi-GAN’s [19] vocoder but modify the original implementation so that it directly accepts the discrete and continuous mixed representation $\mathbf{z}$ as input. The generator includes a set of transposed convolution blocks to increase sample rate of the discrete representation and a residual block with dilated layers to increase the receptive filed. The discriminator contains two types of sub-discriminators: five multi-period discriminators (MPD) and three multi-scale discriminators (MSD).
@@
$L_{G}$ $=\sum_{k=1}^{K}\biggl{[}L_{Adv}(G;D_{k})+\lambda_{fm}L_{FM}(G;D_{k})\biggr{]}$
@@
$L_{Mel}(G)$ $=\mathbb{E}_{x,\hat{x}}\biggl{[}||\phi(x)-\phi(G(z(x)))||_{1}\biggr{]},$ (3)
@@ 3.4 Experimental Setup
The VC experiments are performed among all 90 pairs of 10 test speakers. Each utterance of one test speaker is converted to each of the other 9 speakers’ voices. Each test speaker reads a different set of sentences. All the 10 speakers and their sentences are unseen during training. Therefore, the conversion is non-parallel and zero-shot.
In this experiment, we apply control curves for speed and/or pitch. Four control settings are tested: “No Control” - traditional voice conversion without any explicit control; “Pitch Only” and “Speed Only” denote pitch or speed control but not both; “Speed+Pitch” means both aspects are controlled. We test two curves for pitch control: stressing (i.e., pitch rising abruptly then going down gradually) and rising, and three curves for speed control: parabola, speed up and slow down. The control settings and the control curves are drawn with equal probability for each conversion. We perform both subjective and objective evaluations to assess the conversion quality, intelligibility and controllability of the proposed system.
@@ 3.6 Subjective Evaluation
Audio Quality Test. In the first test, we use mean opinion score (MOS) [35] to assess the naturalness and timbre similarity of the converted speech. Study participants are presented with a set of utterances including one source, one target and several converted utterances. Each set is referred to as a sample. The converted utterances from ControlVC and the two baselines are presented in random order. For each sample, participants are asked to rate between 1-5 on the naturalness of the converted utterances and the timbre similarity between the target and the converted utterance. Higher scores are better. Each participant is asked to complete at least 6 samples, i.e., 36 ratings for 18 converted utterances from 3 comparison methods. Participants are allowed to complete more samples. In total, 233 samples are evaluated, resulting in 1398 ratings. As shown in Fig. 2, ControlVC archives the best MOS among three comparison methods in all control settings. In addition, comparing the three with-control settings with “No Control”, we see that applying controls only slightly decreases the speech quality of the converted speech from ControlVC.
@@
Controllability Test. The second test assesses the controllability of the proposed ControlVC method. We do not include the two baselines in this test due to their poor audio quality in the previous test. The participants are presented with uncontrolled and controlled conversion results, along with a figure of the corresponding control curve(s). The participants are then asked to assess how accurately the curve describes the change of pitch or speed between the uncontrolled and controlled conversions on a scale of 1 to 5, with 1 being “not at all accurate”, 3 being“moderately accurate” and 5 being “very accurate”. Each participant is asked to complete 6 rounds of tests, with each round containing a pitch control, a speed control and a pitch+speed control of the same source-target reference pair. Participants are allowed to complete more rounds. In total, participants completed 169 rounds of tests, resulting in 676 ratings.
@@ 4 Conclusions
In this paper, we proposed a controllable voice conversion system named ControlVC, which allows users to impose time-varying controls on pitch and speed in voice conversion. The converted utterance maintains the source utterance’s linguistic content, mimics the target speaker’s timbre, and sounds natural while following the user input pitch and/or speed controls. Both subjective and objective evaluation results suggest that ControlVC is able to perform pitch and speed control while producing high-quality conversions.